# What Guides Our Choices? Modeling Developers’ Trust and Behavioral Intentions Towards GenAI

### v1

@@ 5.3. Threats to validity and limitations
Construct validity: We captured constructs through self-reported measures, asking participants to express their agreement with literature-derived indicators. This approach assumes that participants’ responses accurately reflect their beliefs and experiences, which might not always be the case due to various biases or misinterpretations. To reduce this threat, we used validated instruments, evaluated the psychometric validity of the PICSE questions, involved practitioners in designing the questions, ran pilot studies, incorporated attention checks, randomized the questionnaire within blocks, and screened the responses. Further, our analysis confirmed that the constructs were internally consistent, reliable, and met convergent and discriminant validity criteria.
@@
Internal validity: Our hypotheses propose associations between constructs, rather than causal relationships, given the cross-sectional nature of the study (Stol and Fitzgerald, 2018). We acknowledge the limitation of self-selection bias, as respondents interested in (or skeptical about) genAI tools might be more willing to complete the questionnaire. Further, a theoretical model like ours cannot capture an exhaustive list of factors. Other factors can play a role, thus positioning our results as a reference for future studies. Future work should also consider using longitudinal data and control for potential confounding factors, such as familiarity with genAI and demographic variables. Additionally, trust is a situation-dependent construct (Jacovi et al., 2021). Although we focused on software development, trust in genAI tools may vary based on finer work contexts (e.g., software design vs. software testing tasks). Therefore, our results should be interpreted as a theoretical starting point, guiding future studies to explore these contextual influences.
@@
External validity: Our survey was conducted within TechNxt and TechCo. While the sample includes engineers from around the globe, it may not fully represent the broader software developer community. However, the sample distribution aligns with previous empirical studies involving software engineers (Russo, 2024; Trinkenreich et al., 2023), providing a suitable starting point to understand the associations presented in our model. The responses were sufficiently consistent to find full or partial empirical support for the hypotheses. Nevertheless, theory development is an iterative process (Shull et al., 2007). Thus, our results should be interpreted as a starting point, aiming for theoretical rather than statistical generalizability. Future studies should replicate, validate, and extend our theoretical model in various contexts.
6. Conclusion
@@
Beyond theoretical contributions, our study offers practical implications, including pointers for design mechanisms that foster appropriate trust and promote equitable user experiences (Sec. 5). While our work enhances the understanding of developers’ trust and behavioral intentions towards genAI tools, long-term longitudinal studies are essential for refining the knowledge of AI adoption dynamics in software engineering.
Acknowledgements.

### v2

@@ V-C Threats to validity and limitations
Construct validity: We captured constructs through self-reported measures, asking participants to express their agreement with literature-derived indicators. This approach assumes that participants’ responses accurately reflect their beliefs and experiences, which might not always be the case due to various biases or misinterpretations. To reduce this threat, we used validated instruments, evaluated the psychometric validity of the PICSE questions, involved practitioners in designing the questions, ran pilot studies, incorporated attention checks, randomized the questionnaire within blocks, and screened the responses. Further, our analysis confirmed that the constructs were internally consistent, reliable, and met convergent and discriminant validity criteria. We did not directly ask participants about their genAI experience; instead, we used their familiarity with genAI tools and its frequency of use at work as proxies to capture their experience with these tools.
@@
Internal validity: Our hypotheses propose associations between constructs, rather than causal relationships, given the cross-sectional nature of the study [102]. We acknowledge the limitation of self-selection bias, as respondents interested in (or skeptical about) genAI tools might be more willing to complete the questionnaire. Further, a theoretical model like ours cannot capture an exhaustive list of factors. Other factors can play a role, thus positioning our results as a reference for future studies. Future work should also consider using longitudinal data and control for potential confounding factors, such as familiarity with genAI and demographic variables. Additionally, trust is a situation-dependent construct [35]. Although we focused on software development, trust in genAI tools may vary based on finer work contexts (e.g., software design vs. software testing tasks). Therefore, our results should be interpreted as a theoretical starting point, guiding future studies to explore these contextual influences.
@@
External validity: Our survey was conducted within GitHub and Microsoft. While the sample includes engineers from around the globe, it may not fully represent the broader software developer community. However, the sample distribution aligns with previous empirical studies involving software engineers [24, 55], providing a suitable starting point to understand the associations presented in our model. The responses were sufficiently consistent to find full or partial empirical support for the hypotheses. Nevertheless, theory development is an iterative process [103]. Thus, our results should be interpreted as a starting point, aiming for theoretical rather than statistical generalizability. Future studies should replicate, validate, and extend our theoretical model in various contexts.
VI Conclusion
@@
Beyond theoretical contributions, our study offers practical implications, including pointers for design mechanisms that foster appropriate trust and promote equitable user experiences (Sec. V). While our work enhances the understanding of developers’ trust and behavioral intentions towards genAI tools, long-term longitudinal studies are essential for refining the knowledge of AI adoption dynamics in software engineering.
Acknowledgments